The one situation where an entire, closed population was tested was the Diamond Princess cruise ship and its quarantine passengers. The case fatality rate there was 1.0%, but this was a largely elderly population, in which the death rate from Covid-19 is much higher. Projecting the Diamond Princess mortality rate onto the age structure of the U.S. population, the death rate among people infected with Covid-19 would be 0.125%. But since this estimate is based on extremely thin data — there were just seven deaths among the 700 infected passengers and crew — the real death rate could stretch from five times lower (0.025%) to five times higher (0.625%). It is also possible that some of the passengers who were infected might die later, and that tourists may have different frequencies of chronic diseases — a risk factor for worse outcomes with SARS-CoV-2 infection — than the general population. Adding these extra sources of uncertainty, reasonable estimates for the case fatality ratio in the general U.S. population vary from 0.05% to 1%.

1) Forgetting for the moment the true fatality rate on Diamond Princess. Assuming we have a sample size of 700 infected persons with a death rate of 0.125%, construct a 99% upper confidence bound on population death rate from Covid-19. (Relevant section is 10.1).

2) The updated situation on Diamond Princess is: 713 infected with 9 fatalities. Compute the sample proportion (fatality rate) and standard error for sample proportion based on these two numbers. (Relevant section is 7.3. Remark: I only understood the number 0.625% in the article after this calculation).

3) Currently about 3.5million infection is reported worldwide with about 250,000 deaths, a fatality rate about 7%. Clearly this rate is much higher than the roughly 1% rate on Diamond Princess. Can you give at least one reason for the difference in these two rates? (This is not a math question, just use your common sense).
Answer #1

Answer! 1 given Sample size n=700 Persons with a death rate P = 0.125%. At 99% confidence interval x=1-0.99 = 0.01 ♡ - 010.1= (0.125-00322), 0:125+0.0322) = (0.0928, 0.1572) The 99%. C. I is = (0.0928, 0.1572) 2) The sample -4 Proportion is given by